OpenStudy (anonymous):

Find the value of the discriminant and describe the number and type of roots. 3x2 - 8x + 2 = 0 Answer 1 x2 + 5x + 6 = 0 Answer 2 6x2 -x - 12 = 0 Answer 3 3x2 - 5x + 6 = 0 Answer 4 x2 - 8x +16 = 0

OpenStudy (raden):

hint : Discriminant (D) = b^2 - 4ac see the values of a,b, and c from the quadratic equ : ax^2 + bx + c = 0

OpenStudy (anonymous):

yeaa..n i honestly need more help cause the power point wont come up

OpenStudy (raden):

1# for you 3x2 - 8x + 2 = 0 here known a = 3, b = -8, and c = 2. therefore the value of D = b^2 - 4ac = (-8)^2 - 4(3)(2) = 64 - 24 = 40 there are 2 real roots notice : if D > 0 then there are 2 real roots if D = 0 then there is one real root if D < 0 then there are 2 imaginary roots

OpenStudy (anonymous):

ok so would it be D=40, two irrational roots??